OpenStudy (anonymous):

Saigon and Shanghai are approximately 2450 miles apart. Two planes leave the two cities and travel towards each other. One plane, travels at 350mph. The other travels at 300mph, but leaves 30 minutes later. How long does it take for the two plane to meet ? ( Based on the first plane ) a) Name a variable for the missing information and what it represents. b) Write an equation between two equivalent expressions. c) Solve the created equation. How long does it takes for the two plane to meet each other? ( Based on the first plane )

OpenStudy (anonymous):

@UnkleRhaukus can you please help me with this ? :(

OpenStudy (anonymous):

Let t is the time: -> d1 = 350t , d2 = 300 ( t - 1/2) Total distance: 350t + 300 ( t - 1/2) = 2450 mi @Rubie Can you solve from here?

OpenStudy (anonymous):

i've done a, just kinnda stuck on b and c

OpenStudy (anonymous):

Don't you see that I already solved b for you ???

OpenStudy (anonymous):

waitt a minute.... so 350t + 300 ( t - 1/2) = 2450 mi is the equation ?

OpenStudy (anonymous):

So you have no idea what the equation looks like?

OpenStudy (anonymous):

" waitt a minute.... so 350t + 300 ( t - 1/2) = 2450 mi is the equation ? " Yes.

OpenStudy (anonymous):

sorry :( I'm new.

OpenStudy (anonymous):

Equation means the 2 sides with the equal "=" sign in the middle!

OpenStudy (anonymous):

okok.

OpenStudy (anonymous):

Take time to read, and think carefully , then ask question to verify your thoughts!

OpenStudy (anonymous):

heyy for c the nest step after the rquation you gave me, is it 350t + ( t -1/2 ) = 2150 ?

OpenStudy (anonymous):

@Rubie Take time to read, and think carefully !!!